# 4DHands: Reconstructing Interactive Hands in 4D with Transformers

### v2

@@ 5 Conclusions
In this paper, we presented 4DHands, a robust solution for recovering interactive hand meshes and their relative movements from monocular inputs. 4DHands is a transformer-based method with novel tokenization and feature fusion strategies. To this end, we introduced the Relation-aware Two-Hand Tokenization (RAT) method, which embeds positional relation information into hand tokens. This allows our network to handle both single-hand and two-hand inputs, leveraging their relative positions to reconstruct detailed hand interactions in real-world scenarios. Moreover, this tokenization approach enhances feature fusion by indicating the relative positions of the hands. To further this goal, we developed a Spatio-temporal Interaction Reasoning (SIR) module that fuses hand tokens in 4D with attention, decoding them into 3D hand meshes and their temporal movements. The effectiveness of 4DHands is demonstrated through extensive evaluations on benchmark datasets, with results from in-the-wild videos and real-world scenarios showcasing its superior performance in interactive hand reconstruction.

### v3

@@ 6 Conclusions
In this paper, we presented OmniHands, a robust solution for recovering interactive hand meshes and their relative movements from monocular inputs. OmniHands is a transformer-based method with novel tokenization and feature fusion strategies. To this end, we introduced the Relation-aware Two-Hand Tokenization (RAT) method, which embeds positional relation information into hand tokens. This allows our network to handle both single-hand and two-hand inputs, leveraging their relative positions to reconstruct detailed hand interactions in real-world scenarios. Moreover, this tokenization approach enhances feature fusion by indicating the relative positions of the hands. To further this goal, we developed a 4D Interaction Reasoning (FIR) module that fuses hand tokens in 4D with attention, decoding them into 3D hand meshes and their temporal movements. The effectiveness of OmniHands is demonstrated through extensive evaluations on benchmark datasets, with results from in-the-wild videos and real-world scenarios showcasing its superior performance in interactive hand reconstruction.